# METANOIA: A Lifelong Intrusion Detection and Investigation System for Mitigating Concept Drift

### v1

@@ Abstract
Index Terms:
Advanced Persistent Threat, Concept Drift, Intrusion Detection, Incremental Learning
@@ I INTRODUCTION
In comparison to raw audit logs, provenance graphs offer a novel perspective for identifying threats and investigating traces, facilitating the differentiation between prolonged normal behavior and malicious activities [1, 2, 3]. Presently, there is a growing trend in leveraging the rich contextual information provided by provenance graphs to perform host-based intrusion detection, referred to as provenance-based intrusion detection systems (PIDSes) [4, 5, 6, 7, 8, 9, 10, 11, 1, 12, 13, 14, 15, 16, 17, 18].
Among them, anomaly-based work defines a model of benign system behavior based on historical log data [10, 11, 1, 12, 13, 14, 15, 16, 17, 18], triggering an alert when the execution significantly deviates from this model. Since it is attack-knowledge-agnostic, counteracts zero-day vulnerabilities, and possesses meticulous detection capabilities, many researchers focus on anomaly-based detection works. Compared to the high latency and expenses associated with offline systems [12, 18, 17, 16], PIDSes capable of processing dynamic evolving provenance graphs in a streaming fashion [1, 8, 14, 15, 13] are better suited for the real-time and cost-effective requirements of current host-based detection.
However, most current anomaly-based PIDSes seem to evade a critical issue: concept drift. Concept drift describes unforeseeable changes in the underlying distribution of streaming data over time [19, 20, 21]. In this work, concept drift is represented as a variation in the benign behavior of a user on the host over time, e.g., switching work environments. This variation can lead to a decline in model performance since the features and patterns learned during training are no longer applicable during testing. In practical scenarios, this common phenomenon corresponds to changes in user behavior at the endpoint, such as switching to different production environments or deploying new software. This change (i.e., concept drift) can result in an unacceptable false-positive rate for existing PIDSes [3, 22, 23].
@@
To feasibly mitigate concept drift (Goal), we introduce incremental learning [24, 25] as a new anomaly detection paradigm (Baseline). Incremental learning refers to the model’s ability to continuously assimilate dynamic data, enabling it to evolve and adapt within an ever-changing environment. Similarly, the system-level streaming logs are dynamic, with their distribution changing over time and continually introducing new entities and interactions. The correspondence between the continuous learning capability of incremental learning and the distribution variability of streaming logs prompts us to consider adopting incremental learning as a new anomaly detection paradigm to mitigate concept drift.
However, directly integrating incremental learning with anomaly-based PIDSes would diminish their intrusion detection capabilities. After thorough research analysis, we identified FOUR CHALLENGES that need to be overcome to integrate incremental learning to mitigate concept drift. First, long-running incremental PIDSes must combat catastrophic forgetting (C1[26]. Furthermore, the PIDSes should endeavor to avoid high false negatives introduced by learning malicious behaviors (C2[27]. Lastly, the PIDSes must achieve precise alerts (C3) and reconstruct attack scenario (C4[15, 14, 3]. It is important to emphasize that these interconnected challenges were progressively uncovered as we attempted to mitigate concept drift. Subsequently, we will elaborate on these challenges.
C1 Combat Catastrophic Forgetting. Some existing studies [12, 15, 14] have adopted the concept of incremental learning, retraining models through oracles (human guidance). However, this approach is plagued by catastrophic forgetting, a phenomenon where models, upon learning new tasks, gradually discard or forget how to perform previous tasks. We argue that anomaly-based PIDSes should be capable of combating catastrophic forgetting, retaining memory of old tasks to reduce redundant training costs and minimize disruptive false alarms.
C2 Avoid Discrimination Paradox. The Rehearsal Approach is a widely acknowledged and effective strategy for mitigating catastrophic forgetting [26]. It fundamentally aids models in retaining the memory of prior learning by revisiting old knowledge. Nonetheless, a specific challenge arises when directly implementing the Rehearsal Approach within endpoint detection: a discrimination paradox. In detail, if the detection system can determine the maliciousness of data selected for replay, subsequent steps (detecting malicious behavior) become redundant. If it fails in this determination, the replay of malicious data might degrade detection performance. Consequently, PIDSes must navigate away from the discrimination paradox.
@@
C4 Reconstruct Attack Scenario. Provenance data can assist in investigating the attack chains leading to intrusions and the potential damage caused by intrusions to the system [2, 28, 29]. However, manually investigating the entire graph is not feasible due to the substantial size of a typical provenance graph and its rapid expansion rate over time [30]. Practical PIDSes need to reconstruct attack scenarios through dependencies among kernel objects, significantly reducing manual work and enabling system administrators to quickly understand intrusions and timely design responses.
@@
  • We evaluate METANOIA on publicly available benchmark datasets from DARPA [31] that simulate APT campaigns, as well as datasets that allow us to fairly compare METANOIA with state-of-the-art PIDSes, UNICORN [1], KAIROS [15], PROGRAPHER [14] and ThreaTrace [13]. Experimental results show that the precision performance of METANOIA on window-level, graph-level, and node-level outperforms existing approaches by 30%, 54%, and 29% respectively.
@@ II-A Scenario
Behavior A and Behavior B are both non-malicious behaviors of the system, and Behavior C is a large-scale APT campaign simulated by DARPA [32]. Since Behavior B and Behavior C are contextually and semantically different from Behavior A, PIDSes recognize them as anomalies and repeat threat alerts.
@@ II-B Prior Approaches & Limitations
Traditional supervised methods, such as ATLAS [33] and APT-KGL [34], require the use of attack samples as training data for manipulation. However, in practice, acquiring such data is extremely challenging. To address this issue, recent PIDSes have adopted unsupervised or semi-supervised detection pipelines. These pipelines model benign data and identify abnormal behaviors as indications of malicious activities. Among state-of-the-art PIDSes, Unicorn [1] and StreamSpot [10] provide graph-level detection (representing the current state of the entire system), ThreaTrace [13] and MAGIC [35] offer entity-level detection, while KAIROS [15] and ShadeWatcher [12] focus on event-level detection.
@@ III Threat Model
We make the following assumptions about the target system. The Trusted Computing Base (TCB) of the system consists of the operating system and audit framework. We assume these components are operating correctly and cannot be manipulated by attackers. We also assume that attackers cannot directly modify the contents of audit logs, which can be achieved through tamper-evident logging techniques [36, 37, 38, 39] beyond the scope of this paper. Finally, we do not consider hardware-level, side-channel, or convert-channel attacks, since kernel-level audit systems do not explicitly capture their behaviors.
It is worth emphasizing that we do not assume the absence of concept drift, which distinguishes our approach from the majority of studies [40, 41, 42, 43, 15, 12, 13, 14]. These studies typically assume that system behaviors remain static over time, i.e., no concept drift exists. However, host behaviors evolve in real-world scenarios with changing user activities and requirements. Therefore, we refrain from making such unrealistic assumptions and instead, METANOIA can serve as an intrusion detection and investigation system for concept drift mitigation using system-level provenance.
@@ IV-B Overview
Preprocess (§ V-A). METANOIA processes streaming data at the level of time windows, which is very common in existing work [15, 14, 44]. METANOIA will perform redundancy filtering of events in the window and encode the events in terms of relation, feature, and state to achieve embedding.
@@
Malice Investigation (§ V-C). METANOIA selects AN with suspicious state (i.e., the results of encoded states in preprocess) as suspicious nodes. Next, METANOIA constructs a mini-graph of suspicious nodes for the current window based on the Steiner Tree algorithm [45, 46] and merges it with the global mini-graph, which is called mini-graph construction. Afterward, METANOIA scores and filters suspicious nodes at the path level to reduce false positives caused by concept drift as much as possible. Finally, METANOIA treats the retained suspicious nodes as malicious nodes, triggers an alert, and reconstructs the attack scenario graph associated with these malicious nodes.
@@ V-A Preprocess
METANOIA constructs a whole-system provenance graph from streaming data collected by logging infrastructures, such as Windows ETW [47], Linux Audit [48], and CamFlow [49]. METANOIA considers three types of kernel objects and eight types of interactions. METANOIA transforms each event into a directed, time-stamped edge, in which the source node represents the event’s subject and the destination node represents the object being acted upon. For any event $e\in E$, METANOIA represents it as a quintuple $\langle{U_{s},U_{o},O,T_{i}}\rangle$. $U_{s}$ and $U_{o}$ are unique identifiers for the subject and object of $e$, respectively. $O$ denotes the type of $e$, and $T_{i}$ denotes the time when $e$ occurred. Table I shows the events between subjects and objects and the node attributes we consider. As shown in Figure 2 (Phase I), METANOIA will perform redundancy filtering and encoding in preprocessing, as described in more detail below.
@@ V-A1 Redundancy Filtering
In order to reduce redundant audit records when detecting and investigating cyber threats, we employed several redundancy filtering techniques from previous works in processing streaming data. Firstly, we implemented the Causality Preserved Reduction (CPR) [50] to aggregate audit records with the same impact scope. For instance, in Figure 1.C, FireFox receives network packets from nhra.com up to 198 times in a single data transmission operation. This is because the operating system typically distributes data among multiple system calls proportionally when performing tasks such as file read/write events. We believe that merging such redundant events does not alter the transmission of information. Another type of redundancy [51] that needs filtering involves events between processes and read-only libraries during initialization. To address this, we established a whitelist of trusted libraries (e.g., ”libc.so”, ”libm.so”) and removed them without compromising the correctness of causal analysis.
@@ V-A2 Encoding
Relation Encoding: METANOIA utilizes one-hot encoding to embed events, similar to most PIDSes [14, 15, 12, 17]. We consider this one-hot encoding approach to be sufficiently simple and effective in assisting the model to differentiate between different types of relations in a flat label space.
Feature Encoding: METANOIA utilizes hierarchical feature hashing [52] to encode node features. Hierarchical feature hashing employs a hash function to map original features into a fixed-size hash space, thereby transforming high-dimensional features into low-dimensional ones. This encoding method is naturally suitable for file paths with hierarchical structures, making it more likely for file names with similar semantics to be mapped to proximate positions in the hash space. For example, the distance between /usr/local/share and /usr/local/lib in feature space will be less than the distance to /usr/ports/shells.
@@ V-B1 Pseudo-Edge Connection
Inspired by the classic network architecture ResNet [53], METANOIA attempts to solve the problem by connecting pseudo-edges for scene replay. There are two key insights of METANOIA: (1) Rehearsal nodes are anomalous yet benign, containing critical information about anomalous scenes, and (2) Pseudo-edge connection assist the model in revisiting anomalous scenes and expanding the scope of recollection. For any event $e$, METANOIA assesses whether a reachable relationship exists between its entity ($U_{s}$ or $U_{o}$) and the rehearsal nodes in the RN Pool. If so, METANOIA establishes a pseudo-edge between them to replay crucial scene information. METANOIA considers that this method helps the model combat catastrophic forgetting through scene recall and prevent the inclusion of malicious scenes. It should be noted that pseudo-edges are only temporarily constructed during the generation of event embedding and do not affect the original components of the provenance graph.
@@ V-B2 Reconstruction Loss Generation
We believe that only by considering all causally related events and their temporal order can one determine whether a node is anomalous. METANOIA employs Temporal Graph Networks (TGN) [54] to reconstruct event types, and generate reconstruction loss (the difference between reconstructed types and actual types).
@@
where $mem$ is a learnable function and METANOIA uses the GRU [55]. Then, METANOIA generates the final embedding $z_{ij}(t)$ for edge $e_{t}$ at any given time $t$:
@@
where $h$ is a learnable function, and METANOIA chooses an attention-based Graph Neural Network (GNN) [56]. $N(i)$ refers to all neighboring nodes of node $i$. $e_{ij}$ refers to the encoding of event $e$ by METANOIA, which is described in detail in Section V-A2. $v_{i}(t)$ refers to the temporal feature of node $i$ at time $t$, which is generated after METANOIA processes the timestamp using a simple linear mapping. In summary, METANOIA embeds the events from node historical information ($m_{i}(t)$), event encoding ($e_{ij}(t)$), and neighborhood structures ($N(i)$).
@@ V-C Malice Investigation
Unlike other existing approaches [14, 12, 1], METANOIA does not directly equate anomaly with maliciousness and issue an alert upon detecting AN but enters a secondary investigation. As shown in Figure 2 (Phase III), METANOIA firstly combines the information of state encoding (Section V-A) to select AN (the output of Phase II) with suspicious states higher than the suspicion threshold $\gamma$ as suspicious nodes. In other words, suspicious nodes are those nodes that simultaneously exhibit suspicious semantics and anomalous behavior. Next, METANOIA obtains the final malicious nodes through mini-graph construction and path-level filtering oriented towards suspicious nodes. The detailed design is as follows:
@@
Typically, there are multiple suspicious nodes in each time window, and METANOIA posits that these suspicious nodes are not independent of each other. The key insight behind mini-graph construction: there are the most anomalous and shortest reachable relationships among suspicious nodes. Based on this insight, METANOIA will construct a mini-tree $MT_{t}$ for all suspicious nodes within the current window $t$ by the Steiner Tree algorithm [45, 46] combined with reconstruction loss. The Steiner Tree algorithm’s primary objective is to locate a tree that encompasses a designated set of specific nodes (i.e., suspicious nodes) in the graph (i.e., provenance graph $G_{t}$ in time window $t$) while minimizing the total weight of the tree as much as possible. Due to the interchangeability between maximum and minimum, METANOIA can utilize this algorithm to find the mini-tree $MT_{t}$, which contains all suspicious nodes within the time window $t$ and has the largest reconstruction loss sum. For a new $MT_{t}$, METANOIA either merges it into the existing mini-graph $MG$ or creates a new mini-graph containing only $MT_{t}$. That is, $MT_{t}$ is merged into the existing $MG$:
@@
0:  Suspicious Nodes $SN$ at Time Window $t$ Mini-Graph $MG$ Path-level Scoring Threshold $\delta$
0:  Malicious Nodes $MN$
1:  for each $sn_{i}$ in $SN$ do
@@
7:     for each $path$ in PATHS do
8:        Volatility     = StandardDeviation($path$)
9:        Periodicity    = 1 / DiscreteFourierTransform($path$)
10:        Heterogeneity = Set(NodeName)/Len($path$)
11:        Persistence    = Len($path$)
@@
14:        if PathScore $>$ $\delta$ then
15:           $MN[sn_{i}]$.append($path$)

### v2

@@ I INTRODUCTION
In comparison to raw audit logs, provenance graphs offer a novel perspective for identifying threats and investigating traces, facilitating the differentiation between prolonged normal behavior and malicious activities [han2020unicorn, nodoze, inam2022sok]. Presently, there is a growing trend in leveraging the rich contextual information provided by provenance graphs to perform host-based intrusion detection, referred to as provenance-based intrusion detection systems (PIDSes) [hossain2017sleuth, morse, zhu2023aptshield, holmes, rapsheet, xiong2020conan, streamspot, FRAPpuccino, han2020unicorn, shadewatcher, threatrace, prographer, kairos, xie2018pagoda, SIGL, wang2020provdetector].
Among them, anomaly-based work defines a model of benign system behavior based on historical log data [streamspot, FRAPpuccino, han2020unicorn, shadewatcher, threatrace, prographer, kairos, xie2018pagoda, SIGL, wang2020provdetector], triggering an alert when the execution significantly deviates from this model. Since it is attack-knowledge-agnostic, counteracts zero-day vulnerabilities, and possesses meticulous detection capabilities, many researchers focus on anomaly-based detection works. Compared to the high latency and expenses associated with offline systems [shadewatcher, wang2020provdetector, SIGL, xie2018pagoda], PIDSes capable of processing dynamic evolving provenance graphs in a streaming fashion [han2020unicorn, rapsheet, prographer, kairos, threatrace] are better suited for the real-time and cost-effective requirements of current host-based detection.
However, most current anomaly-based PIDSes seem to evade a critical issue: concept drift. Concept drift describes unforeseeable changes in the underlying distribution of streaming data over time [tsymbal2004conceptdrift, gama2014conceptdriftsurvey, lu2018conceptdriftlearning]. In this work, concept drift is represented as a variation in the benign behavior of a user on the host over time, e.g., switching work environments. This variation can lead to a decline in model performance since the features and patterns learned during training are no longer applicable during testing. In practical scenarios, this common phenomenon corresponds to changes in user behavior at the endpoint, such as switching to different production environments or deploying new software. This change (i.e., concept drift) can result in an unacceptable false-positive rate for existing PIDSes [inam2022sok, dong2023we, zipperle2022provenance].
@@
To feasibly mitigate concept drift (Goal), we introduce incremental learning [van2022threeincremental, he2011incremental] as a new anomaly detection paradigm (Baseline). Incremental learning refers to the model’s ability to continuously assimilate dynamic data, enabling it to evolve and adapt within an ever-changing environment. Similarly, the system-level streaming logs are dynamic, with their distribution changing over time and continually introducing new entities and interactions. The correspondence between the continuous learning capability of incremental learning and the distribution variability of streaming logs prompts us to consider adopting incremental learning as a new anomaly detection paradigm to mitigate concept drift.
However, directly integrating incremental learning with anomaly-based PIDSes would diminish their intrusion detection capabilities. After thorough research analysis, we identified FOUR CHALLENGES that need to be overcome to integrate incremental learning to mitigate concept drift. First, long-running incremental PIDSes must combat catastrophic forgetting (C1[febrinanto2023lifelongsurvey]. Furthermore, the PIDSes should endeavor to avoid high false negatives introduced by learning malicious behaviors (C2[faber2023lifelongforanomalydetection]. Lastly, the PIDSes must achieve precise alerts (C3) and reconstruct attack scenario (C4[kairos, prographer, inam2022sok]. It is important to emphasize that these interconnected challenges were progressively uncovered as we attempted to mitigate concept drift. Subsequently, we will elaborate on these challenges.
C1 Combat Catastrophic Forgetting. Some existing studies [shadewatcher, kairos, prographer] have adopted the concept of incremental learning, retraining models through oracles (human guidance). However, this approach is plagued by catastrophic forgetting, a phenomenon where models, upon learning new tasks, gradually discard or forget how to perform previous tasks. We argue that anomaly-based PIDSes should be capable of combating catastrophic forgetting, retaining memory of old tasks to reduce redundant training costs and minimize disruptive false alarms.
C2 Avoid Discrimination Paradox. The Rehearsal Approach is a widely acknowledged and effective strategy for mitigating catastrophic forgetting [febrinanto2023lifelongsurvey]. It fundamentally aids models in retaining the memory of prior learning by revisiting old knowledge. Nonetheless, a specific challenge arises when directly implementing the Rehearsal Approach within endpoint detection: a discrimination paradox. In detail, if the detection system can determine the maliciousness of data selected for replay, subsequent steps (detecting malicious behavior) become redundant. If it fails in this determination, the replay of malicious data might degrade detection performance. Consequently, PIDSes must navigate away from the discrimination paradox.
@@
C4 Reconstruct Attack Scenario. Provenance data can assist in investigating the attack chains leading to intrusions and the potential damage caused by intrusions to the system [nodoze, sleuth, king2003backtracking]. However, manually investigating the entire graph is not feasible due to the substantial size of a typical provenance graph and its rapid expansion rate over time [pasquier2017practical]. Practical PIDSes need to reconstruct attack scenarios through dependencies among kernel objects, significantly reducing manual work and enabling system administrators to quickly understand intrusions and timely design responses.
@@
  • We evaluate METANOIA on publicly available benchmark datasets from DARPA [darpa-tc] that simulate APT campaigns, as well as datasets that allow us to fairly compare METANOIA with state-of-the-art PIDSes, UNICORN [han2020unicorn], KAIROS [kairos], PROGRAPHER [prographer] and ThreaTrace [threatrace]. Experimental results show that the precision performance of METANOIA on window-level, graph-level, and node-level outperforms existing approaches by 30%, 54%, and 29% respectively.
@@ II-A Scenario
Behavior A and Behavior B are both non-malicious behaviors of the system, and Behavior C is a large-scale APT campaign simulated by DARPA [darpa5]. Since Behavior B and Behavior C are contextually and semantically different from Behavior A, PIDSes recognize them as anomalies and repeat threat alerts.
@@ II-B Prior Approaches & Limitations
Traditional supervised methods, such as ATLAS [alsaheel2021atlas] and APT-KGL [chen2022aptkgl], require the use of attack samples as training data for manipulation. However, in practice, acquiring such data is extremely challenging. To address this issue, recent PIDSes have adopted unsupervised or semi-supervised detection pipelines. These pipelines model benign data and identify abnormal behaviors as indications of malicious activities. Among state-of-the-art PIDSes, Unicorn [han2020unicorn] and StreamSpot [streamspot] provide graph-level detection (representing the current state of the entire system), ThreaTrace [threatrace] and MAGIC [jia2023magic] offer entity-level detection, while KAIROS [kairos] and ShadeWatcher [shadewatcher] focus on event-level detection.
@@ III Threat Model
We make the following assumptions about the target system. The Trusted Computing Base (TCB) of the system consists of the operating system and audit framework. We assume these components are operating correctly and cannot be manipulated by attackers. We also assume that attackers cannot directly modify the contents of audit logs, which can be achieved through tamper-evident logging techniques [karande2017sgx, paccagnella2020custos, paccagnella2020logging, yagemann2021validating] beyond the scope of this paper. Finally, we do not consider hardware-level, side-channel, or convert-channel attacks, since kernel-level audit systems do not explicitly capture their behaviors.
It is worth emphasizing that we do not assume the absence of concept drift, which distinguishes our approach from the majority of studies [lee2013loggc, ma2016protracer, ma2017mpi, kwon2018mci, kairos, shadewatcher, threatrace, prographer]. These studies typically assume that system behaviors remain static over time, i.e., no concept drift exists. However, host behaviors evolve in real-world scenarios with changing user activities and requirements. Therefore, we refrain from making such unrealistic assumptions and instead, METANOIA can serve as an intrusion detection and investigation system for concept drift mitigation using system-level provenance.
@@ IV-B Overview
Preprocess (§ V-A). METANOIA processes streaming data at the level of time windows, which is very common in existing work [kairos, prographer, li2023nodlink]. METANOIA will perform redundancy filtering of events in the window and encode the events in terms of relation, feature, and state to achieve embedding.
@@
Malice Investigation (§ V-C). METANOIA selects AN with suspicious state (i.e., the results of encoded states in preprocess) as suspicious nodes. Next, METANOIA constructs a mini-graph of suspicious nodes for the current window based on the Steiner Tree algorithm [hwang1992steiner, imase1991dynamicsteiner] and merges it with the global mini-graph, which is called mini-graph construction. Afterward, METANOIA scores and filters suspicious nodes at the path level to reduce false positives caused by concept drift as much as possible. Finally, METANOIA treats the retained suspicious nodes as malicious nodes, triggers an alert, and reconstructs the attack scenario graph associated with these malicious nodes.
@@ V-A Preprocess
METANOIA constructs a whole-system provenance graph from streaming data collected by logging infrastructures, such as Windows ETW [etw], Linux Audit [audit], and CamFlow [pasquier2017camflow]. METANOIA considers three types of kernel objects and eight types of interactions. METANOIA transforms each event into a directed, time-stamped edge, in which the source node represents the event’s subject and the destination node represents the object being acted upon. For any event $e\in E$, METANOIA represents it as a quintuple $\langle{U_{s},U_{o},O,T_{i}}\rangle$. $U_{s}$ and $U_{o}$ are unique identifiers for the subject and object of $e$, respectively. $O$ denotes the type of $e$, and $T_{i}$ denotes the time when $e$ occurred. Table I shows the events between subjects and objects and the node attributes we consider. As shown in Figure 2 (Phase I), METANOIA will perform redundancy filtering and encoding in preprocessing, as described in more detail below.
@@ V-A1 Redundancy Filtering
In order to reduce redundant audit records when detecting and investigating cyber threats, we employed several redundancy filtering techniques from previous works in processing streaming data. Firstly, we implemented the Causality Preserved Reduction (CPR) [xu2016cpr] to aggregate audit records with the same impact scope. For instance, in Figure 1.C, FireFox receives network packets from nhra.com up to 198 times in a single data transmission operation. This is because the operating system typically distributes data among multiple system calls proportionally when performing tasks such as file read/write events. We believe that merging such redundant events does not alter the transmission of information. Another type of redundancy [tang2018nodemerge] that needs filtering involves events between processes and read-only libraries during initialization. To address this, we established a whitelist of trusted libraries (e.g., ”libc.so”, ”libm.so”) and removed them without compromising the correctness of causal analysis.
@@ V-A2 Encoding
Relation Encoding: METANOIA utilizes one-hot encoding to embed events, similar to most PIDSes [prographer, kairos, shadewatcher, SIGL]. We consider this one-hot encoding approach to be sufficiently simple and effective in assisting the model to differentiate between different types of relations in a flat label space.
Feature Encoding: METANOIA utilizes hierarchical feature hashing [zhang2020dynamic] to encode node features. Hierarchical feature hashing employs a hash function to map original features into a fixed-size hash space, thereby transforming high-dimensional features into low-dimensional ones. This encoding method is naturally suitable for file paths with hierarchical structures, making it more likely for file names with similar semantics to be mapped to proximate positions in the hash space. For example, the distance between /usr/local/share and /usr/local/lib in feature space will be less than the distance to /usr/ports/shells.
@@ V-B1 Pseudo-Edge Connection
Inspired by the classic network architecture ResNet [he2016resnet], METANOIA attempts to solve the problem by connecting pseudo-edges for scene replay. There are two key insights of METANOIA: (1) Rehearsal nodes are anomalous yet benign, containing critical information about anomalous scenes, and (2) Pseudo-edge connection assist the model in revisiting anomalous scenes and expanding the scope of recollection. For any event $e$, METANOIA assesses whether a reachable relationship exists between its entity ($U_{s}$ or $U_{o}$) and the rehearsal nodes in the RN Pool. If so, METANOIA establishes a pseudo-edge between them to replay crucial scene information. METANOIA considers that this method helps the model combat catastrophic forgetting through scene recall and prevent the inclusion of malicious scenes. It should be noted that pseudo-edges are only temporarily constructed during the generation of event embedding and do not affect the original components of the provenance graph.
@@ V-B2 Reconstruction Loss Generation
We believe that only by considering all causally related events and their temporal order can one determine whether a node is anomalous. METANOIA employs Temporal Graph Networks (TGN) [rossi2020temporal] to reconstruct event types, and generate reconstruction loss (the difference between reconstructed types and actual types).
@@
where memmem is a learnable function and METANOIA uses the GRU [cho2014gru]. Then, METANOIA generates the final embedding $z_{ij}(t)$ for edge $e_{t}$ at any given time $t$:
@@
where $h$ is a learnable function, and METANOIA chooses an attention-based Graph Neural Network (GNN) [shi2020masked]. $N(i)$ refers to all neighboring nodes of node $i$. $e_{ij}$ refers to the encoding of event $e$ by METANOIA, which is described in detail in Section V-A2. $v_{i}(t)$ refers to the temporal feature of node $i$ at time $t$, which is generated after METANOIA processes the timestamp using a simple linear mapping. In summary, METANOIA embeds the events from node historical information ($m_{i}(t)$), event encoding ($e_{ij}(t)$), and neighborhood structures ($N(i)$).
@@ V-C Malice Investigation
Unlike other existing approaches [prographer, shadewatcher, han2020unicorn], METANOIA does not directly equate anomaly with maliciousness and issue an alert upon detecting AN but enters a secondary investigation. As shown in Figure 2 (Phase III), METANOIA firstly combines the information of state encoding (Section V-A) to select AN (the output of Phase II) with suspicious states higher than the suspicion threshold $\gamma$ as suspicious nodes. In other words, suspicious nodes are those nodes that simultaneously exhibit suspicious semantics and anomalous behavior. Next, METANOIA obtains the final malicious nodes through mini-graph construction and path-level filtering oriented towards suspicious nodes. The detailed design is as follows:
@@
Typically, there are multiple suspicious nodes in each time window, and METANOIA posits that these suspicious nodes are not independent of each other. The key insight behind mini-graph construction: there are the most anomalous and shortest reachable relationships among suspicious nodes. Based on this insight, METANOIA will construct a mini-tree $MT_{t}$ for all suspicious nodes within the current window $t$ by the Steiner Tree algorithm [hwang1992steiner, imase1991dynamicsteiner] combined with reconstruction loss. The Steiner Tree algorithm’s primary objective is to locate a tree that encompasses a designated set of specific nodes (i.e., suspicious nodes) in the graph (i.e., provenance graph $G_{t}$ in time window $t$) while minimizing the total weight of the tree as much as possible. Due to the interchangeability between maximum and minimum, METANOIA can utilize this algorithm to find the mini-tree $MT_{t}$, which contains all suspicious nodes within the time window $t$ and has the largest reconstruction loss sum. For a new $MT_{t}$, METANOIA either merges it into the existing mini-graph $MG$ or creates a new mini-graph containing only $MT_{t}$. That is, $MT_{t}$ is merged into the existing $MG$:
@@
0: Suspicious Nodes SNSN at Time Window $t$ Mini-Graph MGMG Path-level Scoring Threshold $\delta$
0: Malicious Nodes MNMN
1:for each $sn_{i}$ in SNSN do
@@
7:  for each pathpath in PATHS do
8:   Volatility = StandardDeviation(pathpath)
9:   Periodicity = 1 / DiscreteFourierTransform(pathpath)
10:   Heterogeneity = Set(NodeName)/Len(pathpath)
11:   Persistence = Len(pathpath)
@@
14:   if PathScore >> $\delta$ then
15:    $MN[sn_{i}]$.append(pathpath)